OpenStudy (anonymous):

Which of the following does not make sense for the series : 2+4+8+16+32+64+128

OpenStudy (anonymous):

\[\sum_{n=1}^{7} 2^2 or \sum_{n=0}^{6} 2^n+1 or \sum_{n=2}^{8} 128 (1/2)^ {n-2} OR \sum_{n=1}^{7} 128 (1/2)^{n} \]

OpenStudy (anonymous):

(( actually on the first series it has to be 2^n not 2^2)

ganeshie8 (ganeshie8):

you figured that out correctly ! you still stuck wid something here ?

OpenStudy (anonymous):

Wait no I didn't, I just did a type!

OpenStudy (anonymous):

*typo!

ganeshie8 (ganeshie8):

for second option, that +1 is at the exponent, right ?

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

good :) then except first option, all others are making sense to me. what about u ?

OpenStudy (anonymous):

But one of them is incorrect??

ganeshie8 (ganeshie8):

first one doesn't make sense

OpenStudy (anonymous):

It is supposed to be 2^n because I did a typo

ganeshie8 (ganeshie8):

oh i see ok

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

254

ganeshie8 (ganeshie8):

\(\sum_{n=1}^{7} 2^n \ or\ \sum_{n=0}^{6} 2^n+1 \ or\ \sum_{n=2}^{8} 128 (1/2)^ {n-2} \ OR \ \sum_{n=1}^{7} 128 (1/2)^{n} \)

OpenStudy (anonymous):

the secone one is actually ^n+1

ganeshie8 (ganeshie8):

\(\sum_{n=1}^{7} 2^n \ or\ \sum_{n=0}^{6} 2^{n+1} \ or\ \sum_{n=2}^{8} 128 (1/2)^ {n-2} \ OR \ \sum_{n=1}^{7} 128 (1/2)^{n} \)

OpenStudy (anonymous):

Yes!

ganeshie8 (ganeshie8):

First option works

ganeshie8 (ganeshie8):

second option works

ganeshie8 (ganeshie8):

third option works

OpenStudy (anonymous):

oh but the 4th doesn't

OpenStudy (anonymous):

but doesn't the 3rd not work aswell

ganeshie8 (ganeshie8):

fourth option if u expand, lets see what we get :- \( \sum_{n=1}^{7} 128 (1/2)^{n} \) \( 128 (1/2)^{1} + 128 (1/2)^{2} + 128 (1/2)^{3} + 128 (1/2)^{4} + 128 (1/2)^{5} + 128 (1/2)^{6} + 128 (1/2)^{7} \) \( 64+ 32 + 16 + 8 + 4 + 2 + 1 \)

OpenStudy (anonymous):

Oh h I see

ganeshie8 (ganeshie8):

so fourth option wont work.

ganeshie8 (ganeshie8):

lets expand third option and see

ganeshie8 (ganeshie8):

\(\sum_{n=2}^{8} 128 (1/2)^ {n-2} \) \( 128 (1/2)^{2-2} + 128 (1/2)^{3-2} + 128 (1/2)^{4-2} + 128 (1/2)^{5-2} + 128 (1/2)^{6-2} + 128 (1/2)^{7-2} + 128 (1/2)^{8-2} \) \( 128 (1/2)^{0} + 128 (1/2)^{1} + 128 (1/2)^{2} + 128 (1/2)^{3} + 128 (1/2)^{4} + 128 (1/2)^{5} + 128 (1/2)^{6} \) \( 128 + 64 + 32 + 16 + 8 + 4 + 2 \)

OpenStudy (anonymous):

mhm so that would work ... ?

ganeshie8 (ganeshie8):

yep ! only fourth doesnt work, rest everything works

OpenStudy (anonymous):

But how come it still works but it is just flipped ?

OpenStudy (anonymous):

Also wheny ou plug in "n" for #3 it seems to not make sense ?

ganeshie8 (ganeshie8):

128 + 64 + 32 + 16 + 8 + 4 + 2 is same as 2 + 4 + 8 + 16 + 32 + 64 + 128

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

I did not know that lol

ganeshie8 (ganeshie8):

lol

OpenStudy (anonymous):

thnx!